OpenStudy (anonymous):

PLEASE HELP. SIMPLE INTEREST PROBLEM. A woman invested $8000 into mutual funds in one year part of your investment earned 7% simple interest and the other part earned 6% simple interest. At the end of the year the woman had earned $512. how much did she invest at each rate?

OpenStudy (anonymous):

i got .07x+.06y=512

OpenStudy (anonymous):

and x+y=8000

OpenStudy (anonymous):

She invested x amount with 6% and y amount with 7% We know she had 8000 cash, so x+y=8000 The interest with 6% is 0.06 that is 0.06x , same for y 0.07y They totaled a profit of 512 thus 0.06x+0.07y=512

OpenStudy (anonymous):

i just am not sure what to do from there

OpenStudy (anonymous):

That is good :)

OpenStudy (anonymous):

now i use substitution or elimination? idk how

OpenStudy (anonymous):

Well just solve them :) It looks a bit nasty

OpenStudy (anonymous):

I DONT KNOW HOW TO SOLVE THEM

OpenStudy (anonymous):

x+y=8000 .07x+.06y=512 Multiply by 100 to simplify 7x+6y=51200 First equation multiply by 6 6x+6y=48000 Take away second from first 7x+6y=51200 minus 6x+6y=48000 x=3200 Do not yell at someone ever!! Especially when they are helping you..... really rude

OpenStudy (anonymous):

lmao i put it in caps by accident I'm sorry! I'm not yelling at you!! i appreciate your help, honestly.

OpenStudy (anonymous):

OK, never mind. This was the method of elimination. I made 6y for both equations so that I could take away one from the other

OpenStudy (anonymous):

thankyou so much sir i understand now